OpenStudy (anonymous):

Use the Law of Cosines to find the missing angle. Find m∠A to the nearest tenth of a degree.

OpenStudy (anonymous):



OpenStudy (anonymous):

answer choice 33.9º 57.7º 46.3º 85.7º

OpenStudy (anonymous):

a^2=b^2+c^2-2bc*cosA 〖17〗^2=〖22〗^2+〖30〗^2-2*22*30*cosA cosA=0.8295 =>A=33.94